OpenStudy (amenah8):

Medicine capsules are shaped like cylinders with hemispherical ends. Suppose someone wants to put a dose of 1ml of a drug into the capsule. What should the dimensions of the capsule be that use the least amount of material to contain the dose?

OpenStudy (amenah8):

V cylinder = pi * r^2 * h V hemisphere = 2/3 * pi * r

OpenStudy (amenah8):